OpenStudy (anonymous):

Solve for x: -2(x - 3) = 10 A x = -2 B x = 2 C x = -8 D x = 8

OpenStudy (mathstudent55):

First distribute on left side. -2(x - 3) = 10 -2x + 6 = 10 Now subtract 6 from both sides: -2x = 4 Now divide both sides by -2 x = -2